OpenStudy (anonymous):

The area of an isosceles triangle is 60. Length of each equal side is 13 find base?

OpenStudy (anonymous):



OpenStudy (anonymous):

x^2+y^2 = 169 xy=60 x=60/y (60/y)^2+y^2 = 169

OpenStudy (anonymous):

If you solve the last equation for y then multiply by 2 you will get the base.